OpenStudy (anonymous):

Evaluate log^5 1/625 a.-3 b.5 c.-4 d.4

OpenStudy (anonymous):

Is it: \(\LARGE \log_{5}{\frac{1}{625}} \)? If so, think about which number 5 needs to be raised to to get 625. In other words: 5^x = 625, what is x?

OpenStudy (anonymous):

yea thats it and is it 4

OpenStudy (anonymous):

It's 4 for 5^x = 625. Now, think about it this way: we want a number such that 5^x = 1/625. Remember that 5^(-x) = 1/5^x and we have that 5^4 = 625, so the answer should be?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yup :-)